OpenStudy (anonymous):

(14 - a^0b^2)^-3 for a = -2 and b = 4

jimthompson5910 (jim_thompson5910):

(14 - a^0b^2)^(-3) (14 - (-2)^0*(4)^2)^(-3) (14 - (1)*(16))^(-3) (14 - 16)^(-3) (-2)^(-3) 1/((-2)^3) 1/(-8) -1/8 So (14 - a^0b^2)^(-3) = -1/8 when a = -2 and b = 4.